OpenStudy (goldenwolf13):

Please help (3.8*10^4)(2.5*10^4)

OpenStudy (kaiiskool):

First thing you want to do is exponents

RhondaSommer (rhondasommer):

\[\Huge\color{#3debf5}{\mathbb{WELCOME\\TO\\OPENSTUDY}}\]

OpenStudy (abdullahm):

Here is an example: (5.2 * 10^3)(3.2 * 10^3) 5.2 * 3.2 = 16.64 \(10^3 * 10^3 = 10^{3+3} = 10^6\) And so we get: \((5.2 * 10^3)(3.2 * 10^3) = 16.64 \times 10^6 = 1.664 \times 10^7\)

OpenStudy (goldenwolf13):

Thank you@AbdullahM

OpenStudy (abdullahm):

@goldenwolf13 First do 3.8 * 2.5 What is that equal to? Now, what is 10^4 * 10^4 equal to? When multiplying two numbers with the same base, in this case 10, you have to add the exponents!

OpenStudy (goldenwolf13):

Can i ask u one more question

OpenStudy (abdullahm):

Go ahead!

OpenStudy (goldenwolf13):

Ok. Whats 8.33*10^3 __________ 5.4*10^5 Simplifyed

OpenStudy (abdullahm):

This time we have to simply divide 8.33 by 5.4 And then we divide 10^3 by 10^5 When we divide exponents with the same base, we have to subtract the exponents. For example, \(\Large\bf \frac{10^6}{10^9} = 10^{6-9} =10^{-3}\)

OpenStudy (goldenwolf13):

Ok thank you @AbdullahM awesome name by the way

OpenStudy (abdullahm):

It was my pleasure! And thanks for the compliment!

OpenStudy (goldenwolf13):

No problem